OpenStudy (anonymous):

I need help solvng this sytem of equation and determining whether or not there is a solution(if so what is the solution? x+7y=2, x=4-7y

OpenStudy (anonymous):

plug x=4-7y in the first equation... (4-7y) + 7y =2 and then solve for y. then plug that answer into the second equation to get x. Does that help?

OpenStudy (anonymous):

somewhat but I still dont know how determine if it is a solution or not

OpenStudy (anonymous):

Sorry..didn't work that all the way out. When you solve it, you will find that the x or y will cancel out depending on the equation you use. These are two parallel lines so there is no (x,y) point that will make both equations true. All you can have is y= (2-x)/7 and x = 4-7y

OpenStudy (anonymous):

so there is no solution

OpenStudy (anonymous):

Correct. no solution.

OpenStudy (anonymous):

thanks